x=1/2y+3
2x-y=6
please help solve these linear equations by using substitution.

Answer by RAY100(1637)

This is a trick question
1)x=1/2y+3
multiply by 2
2x=y+6
or 2x-y=6
this is exactly what the second equation is therefore
the lines coincide, lay on top of each other,
and have infinite solutions.
The other case to watch for is two lines with the same slope, but different y intersects
for example 1) x+y=5, and 2) x+y= 10
These have no solution, the lines do not intersect